OpenStudy (anonymous):

A rectangle is inscribed with its base on the x-axis and its upper corners on the parabola y=4−x2. What are the dimensions of such a rectangle with the greatest possible area?

OpenStudy (mertsj):

2sqrt3/3 by 8/3

OpenStudy (anonymous):

l=2*(4/3)^0.5 h=4-(4/3)^2=8/3

OpenStudy (anonymous):

myininaya can u come help me when u r finished?

myininaya (myininaya):

 base is a-(-a)=a+a=2a height is y=4-a^2 Area=base*height=2a(4-a^2) Area=8a-2a^3 A'=8-6a^2 set A'=0 8-6a^2=0 solve for a 8=6a^2 divide 6 on both sides 8/6=a^2 4/3=a^2 so \[a=\pm \sqrt{\frac{4}{3}}=\pm \frac{2}{\sqrt{3}} \cdot \frac{\sqrt{3}}{\sqrt{3}}=\pm \frac{2 \sqrt{3}}{3}\] so we have \[a=\frac{2 \sqrt{3}}{3}\]